OpenStudy (anonymous):

Help please? Tell whether the given equation has the ordered pair as a solution? y= x + 2/3? (1, 1/3) I know how to do this, I am just confused on this problem..

OpenStudy (luigi0210):

Plug them in, the points are usually (x, y)

OpenStudy (anonymous):

I know, I did its 1/3 = 1 + 2/3

OpenStudy (luigi0210):

If the statement is true, then the equation is true.

OpenStudy (anonymous):

I think its not true?

OpenStudy (luigi0210):

does 1/3=5/3?

OpenStudy (anonymous):

No

OpenStudy (luigi0210):

The it's false

OpenStudy (anonymous):

Thats what I was thinking, Thank you once again :)

OpenStudy (luigi0210):

Anytime :)